OpenStudy (anonymous):

find the measure of an angle whose supplement is 3 times its complement?

Parth (parthkohli):

\[\text{Compliment of an angle} \\ \\ \quad \text{the angle measure which added to the original angle results in 90}^\circ\].For example, the compliment of 67 is 23.

Parth (parthkohli):

Similarly, two supplementary angles add up to get 180 degrees.

Parth (parthkohli):

So, the supplement of an angle (suppose theta) is \(180 - \theta\). The complement of theta is \(9 0 -\theta \).

Parth (parthkohli):

We have:\[ 180-\theta = 3(90 - \theta)\]

Parth (parthkohli):

Solve for theta, and it's just like solving for a variable.

Parth (parthkohli):

@Faye281 Rest for you